# Characterizing the support of semiclassical measures for higher-dimensional cat maps

### v1

@@ 2.2. Symbol calculus
These derivative bounds naturally induce a family of seminorms of $S_{L,\rho,\rho^{\prime}}$. We now quote the following properties of $S_{L,\rho,\rho^{\prime}}$.
@@
    $\operatorname{Op}_{h}(a)\operatorname{Op}_{h}(b)=\operatorname{Op}_{h}(a\#b)$, where $a\#b$ satisfies the following asymptotic expansion as $h\rightarrow 0$ for all $l\in\mathbb{N}$:
@@ 2.3. Symbols on torus
    $\operatorname{Op}_{N,\theta}(a)\operatorname{Op}_{N,\theta}(b)=\operatorname{Op}_{N,\theta}(a\#b)$ where $a\#b$ satisfies the following asymptotic expansion for all $l\in\mathbb{N}$:
@@
    $\|\operatorname{Op}_{N,\theta}(a)\|_{\mathcal{H}_{N}(\theta)\rightarrow\mathcal{H}_{N}(\theta)}$ is bounded by some $S_{L,\rho,\rho^{\prime}}$-seminorm of $a$.
@@
    $\left\langle\operatorname{Op}_{N,\theta}(a)f,f\right\rangle_{\mathcal{H}_{N}(\theta)}\geq-CN^{\rho+\rho^{\prime}-1}\|f\|_{\mathcal{H}_{N}(\theta)}^{2}\quad\text{for all}\quad f\in\mathcal{H}_{N}(\theta).$
@@
When $\rho=\rho^{\prime}<1/2$, $S_{L,\rho,\rho^{\prime}}(\mathbb{R}^{2n})$ becomes the following standard symbol class, see e.g. [z12semiclassical]*§4.4:
@@ 2.4.1. Basic uncertainty principle
Fix $\chi\in C_{c}^{\infty}(\mathbb{R}^{d})$, $\delta\in[0,1]$, and $x^{\prime}_{0},\xi^{\prime}_{0}\in\mathbb{R}^{d}$. Then
@@
Let $L_{a}$ and $L_{b}$ be coisotropic subspaces such that $L_{a}\subset\{x^{\prime}=0\}$ and $L_{b}\subset\{\xi^{\prime}=0\}$. Fix $x^{\prime}_{0},\xi^{\prime}_{0}\in\mathbb{R}^{d}$ and suppose $a\in S_{L_{a},\rho+\varepsilon,\varepsilon}(\mathbb{R}^{2n})$ and $b\in S_{L_{b},\rho+\varepsilon,\varepsilon}(\mathbb{R}^{2n})$ with $\operatorname{supp}a\subset\{|x^{\prime}-x^{\prime}_{0}|\leq h^{\delta}\}$ and $\operatorname{supp}b\subset\{|\xi^{\prime}-\xi^{\prime}_{0}|\leq h^{\delta}\}$ for some $\delta\leq\rho+\varepsilon$. Let $\delta^{\prime}=\min\{1-\rho-2\varepsilon,d(2\delta-1)/2\}$. Then
@@
[…]
@@ 2.4.2. Higher-dimensional fractal uncertainty principle
We begin with some definitions.
@@
[…]
@@
(2.20) $S(\delta)\coloneqq S+B_{\delta}(0),$
where we use the notation $B_{r}(x)$ to denote the ball centered at $x$ of radius $r$.
@@
(2.21) $\sup|\partial^{\alpha}\chi_{\pm}|\leq C_{\alpha}h^{-|\alpha|}.$
Therefore,
@@
Note that $\chi_{j}^{\pm}$ still satisfy the derivative bounds (2.21) for $C_{\alpha}$ depending only on $\alpha$. With respect to the strong operator topology,
@@ 2. We will show
[…]
@@
[…]
@@ 3.1. Definitions
with corresponding symbol
@@
We call elements of $\mathcal{X}$ uncontrolled long logarithmic words and elements of $\mathcal{Y}$ controlled long logarithmic words. To simply notation, we set
@@ 3.2. Decay of $B_{\mathcal{Y}}$
[…]
@@
$=N^{-\frac{2\rho\log 2}{J\log\Lambda}}\left(\left\langle\operatorname{Op}_{N,\theta}\left(b_{d}\right)^{*}\operatorname{Op}_{N,\theta}\left(b_{d}\right)u,u\right\rangle_{\mathcal{H}}-\left\langle\operatorname{Op}_{N,\theta}\left(b_{c}\right)^{*}\operatorname{Op}_{N,\theta}\left(b_{c}\right)u,u\right\rangle_{\mathcal{H}}\right)+CN^{-\frac{4}{5}}\|u\|^{2}_{\mathcal{H}}$
@@
$=N^{-\frac{2\rho\log 2}{J\log\Lambda}}\left(\left\|B_{d}u\right\|^{2}_{\mathcal{H}}-\left\|B_{c}u\right\|^{2}_{\mathcal{H}}\right)+CN^{-\frac{4}{5}}\|u\|^{2}_{\mathcal{H}}.$
@@ 3.3. Decay of $B_{\mathcal{X}}$
#$\#\mathcal{X}\leq C(\log N)^{2J}N^{\frac{2c\rho\sqrt{\alpha}}{\log\Lambda}}.$
@@
As #$\#\mathcal{X}=\#(\mathcal{W}(T_{0})\setminus\mathcal{Z})^{2J}$, it suffices to bound #$\#(\mathcal{W}(T_{0})\setminus\mathcal{Z})$. Recall that $\mathrm{w}=w_{0}\cdots w_{T_{0}-1}$ is in $\mathcal{W}(T_{0})\setminus\mathcal{Z}$ if and only if #$\#\{j\in\{0,\ldots,T_{0}-1\}:w_{j}=1\}<\alpha T_{0}$. Thus, using the fact that $\alpha<1/2$ and $T_{0}=\left\lfloor\frac{\rho\log N}{J\log\Lambda}\right\rfloor$,
#$\#(\mathcal{W}(T_{0})\setminus\mathcal{Z})$ $\leq\sum_{0\leq l\leq\alpha T_{0}}\binom{T_{0}}{l}\leq(\alpha T_{0}+1)\binom{T_{0}}{\left\lceil\alpha T_{0}\right\rceil}$
@@
$-(\alpha\log\alpha+(1-\alpha)\log(1-\alpha))\leq c\sqrt{\alpha},$
@@ 4.1. Proof groundwork
We fix a hyperbolic $A\in\operatorname{Sp}(2n,\mathbb{Z})$ that is diagonalizable in $\mathbb{C}$. Let $M\in\mathcal{M}_{A}$ be a metaplectic transformation quantizing $A$. As $A$ is symplectic, its transpose is conjugate to its inverse. Thus, if $\lambda$ is an eigenvalue of $A$, $\lambda^{-1}$ is also an eigenvalue.
@@ 4.1.3. Preliminary bound on $B_{\mathrm{w}}$
We use the notation $\mathrm{z}_{(2i-1)}\coloneqq z_{0}^{(2i-1)}\cdots z_{s_{i}-1}^{(2i-1)}$ and $\mathrm{z}_{(2i)}=z_{0}^{(2i)}\cdots z_{r_{i}-1}^{(2i)}$ and set
$\tilde{b}_{\mathrm{z}_{(2i-1)}}=\prod_{k=0}^{s_{i}-1}\tilde{b}_{z_{k}^{(2i-1)}}\circ A^{k}.$
@@
$\tilde{b}_{\mathrm{z}_{m}}=\prod_{k=0}^{2T_{1}-1}\tilde{b}_{z_{k}}\circ A^{k},$
@@
$\tilde{B}_{\mathrm{z}_{m}}\coloneqq\tilde{B}_{z_{2T_{1}-1}}(2T_{1}-1)\cdots\tilde{B}_{z_{1}}(1)\tilde{B}_{z_{0}}(0),$
@@
$\left\|\sum_{\mathrm{z}_{m}\in\mathbb{M}_{2}}\tilde{B}_{\mathrm{z}_{m}}\right\|\leq SN^{\frac{2\alpha\rho\log SR}{\log\Lambda}}\max_{\mathrm{z}_{m}\in\mathbb{M}}\|B_{\mathrm{z}_{m}}\|.$
@@
It suffices to show #$\#\mathbb{M}_{2}\leq SN^{\frac{2\alpha\rho\log SR}{\log\Lambda}}$. We fix $11\leq i\leq K+$ 1 and examine $\tilde{b}_{\mathrm{z}_{(2i-1)}}$. Note that for each $1\leq k\leq S$, there exists a unique $1\leq j_{k}\leq S$ such that $A(\pi_{W^{\perp\sigma}}(x_{k}))=\pi_{W^{\perp\sigma}}(x_{j_{k}})$. In other words, $A(\operatorname{supp}\tilde{\varphi}_{R+k})\cap\operatorname{supp}\tilde{\varphi}_{R+j_{k}}\neq\emptyset$. We know
@@
$z_{1}^{(2i-1)}=j_{z_{0}^{(2i-1)}},\quad z_{2}^{(2i-1)}=j_{z_{1}^{(2i-1)}},\quad\ldots,\quad z_{s_{i}-1}^{(2i-1)}=j_{z_{s_{i}-2}^{(2i-1)}}.$
Consequently, for each choice of $z_{0}^{(2i-1)}\in\{R+1,\ldots,R+S\}$, there is only one way to pick $z_{1}^{(2i-1)},\ldots,z_{s_{i}-1}^{(2i-1)}$ such that $\tilde{b}_{\mathrm{z}_{(2i-1)}}$ is nonzero. Then, there are $S$ nonzero choices of each $\tilde{b}_{\mathrm{z}_{(2i-1)}}$. Therefore, #$\#\mathbb{M}_{2}\leq S^{K+1}R^{2\alpha T_{1}}$. By (3.3) and (4.9),
@@
We now turn our attention to $\sum_{\mathrm{z}_{m}\in\mathbb{M}_{1}}\tilde{B}_{\mathrm{z}_{m}}$
@@
Choose $\mathrm{z}_{m}\in\mathbb{M}_{1}$. From the proof of Lemma 4.2, we see that as $\mathrm{z}_{m}\in\mathbb{M}_{1}$, there exists some $10\leq k<2T_{1}-$ 1 such that
$\operatorname{supp}(\tilde{b}_{z_{k}}\circ A^{k})\cap\operatorname{supp}(\tilde{b}_{z_{k+1}}\circ A^{k+1})=\emptyset.$
@@
$\|\tilde{B}_{z_{k+1}}(k+1)\tilde{B}_{z_{k}}(k)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}=\mathcal{O}(N_{j}^{-\infty}).$
@@
where $h=(2\pi N)^{-1}$ and $\operatorname{Op}_{h}$ is the semiclassical Weyl quantization, defined in (2.1). Thus, using the exact Egorov’s Theorem (2.6) and the fact that $M$ is unitary,
$\|\tilde{B}_{z_{i}}(i)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\leq\|\operatorname{Op}_{h_{j}}(\tilde{b}_{z_{i}}\circ A^{i})\|_{L^{2}\rightarrow L^{2}}=\|\operatorname{Op}_{h_{j}}(\tilde{b}_{z_{i}})\|_{L^{2}\rightarrow L^{2}}.$
@@
$\|\tilde{B}_{\mathrm{z}_{m}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}$ $\leq\|\tilde{B}_{z_{2T_{1}-1}}(2T_{1}-1)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\cdots\|\tilde{B}_{z_{k+1}}(k+1)\tilde{B}_{z_{k}}(k)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\cdots\|\tilde{B}_{z_{0}}(0)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}$
@@
(4.11) $\|B_{\mathrm{w}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\leq SN^{\frac{2\alpha\rho\log SR}{\log\Lambda}}\max_{\mathrm{z}_{m}\in\mathbb{M}}\|\tilde{B}_{\mathrm{z}_{m}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}+\mathcal{O}(N_{j}^{-\infty}).$
@@
Write $\mathrm{z}_{m}=\mathrm{z}_{+}\mathrm{z}_{-}$, where $\mathrm{z}_{\pm}$ are words of length $T_{1}$. Relabel $\mathrm{z}_{+}=z_{T_{1}}^{+}\cdots z_{1}^{+}$, $\mathrm{z}_{-}=z_{0}^{-}\cdots z_{T_{1}-1}^{-}$. We see that
(4.14) $\tilde{B}_{\mathrm{z}_{m}}$ $=\tilde{B}_{z^{-}_{T_{1}-1}}(2T_{1}-1)\cdots\tilde{B}_{z^{-}_{0}}(T_{1})\tilde{B}_{z_{1}^{+}}(T_{1}-1)\cdots\tilde{B}_{z_{T_{1}-1}^{+}}(0)$
@@
$\tilde{b}_{+}=\prod_{k=1}^{T_{1}}\tilde{b}_{z_{k}^{+}}\circ A^{-k}\quad\text{and}\quad\tilde{b}_{-}=\prod_{k=0}^{T_{1}-1}\tilde{b}_{z_{k}^{-}}\circ A^{k}.$
@@
We give the proof for $\tilde{b}_{-}$; the proof for $\tilde{b}_{+}$ follows similarly. By Lemma 2.8, it suffices to show for $i=1,\ldots,R+S$ and $10\leq k\leq T_{1}-$ 1, each $S_{L_{-},\rho,0}$-seminorm of $\tilde{b}_{i}\circ A^{k}$ is bounded uniformly in $k$ and $N$.

### v2

@@ 2.2. Symbol calculus
These derivative bounds naturally induce a family of seminorms of $S_{L,\rho,\rho^{\prime}}$.
When $\rho=\rho^{\prime}<1/2$, $S_{L,\rho,\rho^{\prime}}(\mathbb{R}^{2n})$ becomes the following symbol class, see e.g. [z12semiclassical]*§4.4:
We now quote the following properties of $S_{L,\rho,\rho^{\prime}}$.
@@
    $\operatorname{Op}_{h}(a)\operatorname{Op}_{h}(b)=\operatorname{Op}_{h}(a\#b)$, where a#ba\#b satisfies the following asymptotic expansion as $h\rightarrow 0$ for all $l\in\mathbb{N}$:
@@ 2.3. Symbols on torus
    $\operatorname{Op}_{N,\theta}(a)\operatorname{Op}_{N,\theta}(b)=\operatorname{Op}_{N,\theta}(a\#b)$ where a#ba\#b satisfies the following asymptotic expansion for all $l\in\mathbb{N}$:
@@
    $\|\operatorname{Op}_{N,\theta}(a)\|_{\mathcal{H}_{N}(\theta)\rightarrow\mathcal{H}_{N}(\theta)}$ is bounded by some $S_{L,\rho,\rho^{\prime}}$-seminorm of $a$, where the choice of the seminorm depends only on $n,\rho$ and $\rho^{\prime}$.
@@
    $\left\langle\operatorname{Op}_{N,\theta}(a)f,f\right\rangle_{\mathcal{H}_{N}(\theta)}\geq-C_{a}N^{\rho+\rho^{\prime}-1}\|f\|_{\mathcal{H}_{N}(\theta)}^{2}\quad\text{for all}\quad f\in\mathcal{H}_{N}(\theta).$
@@ 2.4.1. Basic uncertainty principle
Fix $\chi\in C_{c}^{\infty}(\mathbb{R}^{d})$, $\delta\in[\tfrac{1}{2},1]$, and $x^{\prime}_{0},\xi^{\prime}_{0}\in\mathbb{R}^{d}$. Then
@@
Let $L_{a}$ and $L_{b}$ be coisotropic subspaces such that $L_{a}\subset\{x^{\prime}=0\}$ and $L_{b}\subset\{\xi^{\prime}=0\}$. Fix $x^{\prime}_{0},\xi^{\prime}_{0}\in\mathbb{R}^{d}$ and suppose $a\in S_{L_{a},\rho+\varepsilon,\varepsilon}(\mathbb{R}^{2n})$ and $b\in S_{L_{b},\rho+\varepsilon,\varepsilon}(\mathbb{R}^{2n})$ with $\operatorname{supp}a\subset\{|x^{\prime}-x^{\prime}_{0}|\leq h^{\delta}\}$ and $\operatorname{supp}b\subset\{|\xi^{\prime}-\xi^{\prime}_{0}|\leq h^{\delta}\}$, where $\rho\in(\tfrac{1}{2},1)$ with $\rho+2\varepsilon<1$ and $\delta\in(\tfrac{1}{2},\rho+\varepsilon]$. Let $\delta^{\prime}=\min\{1-\rho-2\varepsilon,d(2\delta-1)/2\}$. Then
@@
We use $\|\cdot\|$ to denote $\|\cdot\|_{L^{2}(\mathbb{R}^{n})\rightarrow L^{2}(\mathbb{R}^{n})}$. Let $\tilde{\chi}\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\tilde{\chi}=1$ on $\{|x^{\prime}|\leq 1\}$ and $\operatorname{supp}\tilde{\chi}\subset\{|x^{\prime}|\leq 2\}$. Define $\chi_{x},\chi_{\xi}\in C^{\infty}(\mathbb{R}^{2n})$ by $\chi_{x}(x,\xi)=\tilde{\chi}(\frac{x^{\prime}-x_{0}^{\prime}}{h^{\delta}})$ and $\chi_{\xi}(x,\xi)=\tilde{\chi}(\frac{\xi^{\prime}-\xi^{\prime}_{0}}{h^{\delta}})$. Since $L_{a}\subset\{x^{\prime}=0\}$, $L_{b}\subset\{\xi^{\prime}=0\}$, and $\delta\leq\rho+\varepsilon$, we know $\chi_{x}\in S_{L_{a},\rho+\varepsilon,\varepsilon}$ and $\chi_{\xi}\in S_{L_{b},\rho+\varepsilon,\varepsilon}$. Then using Lemma 2.6 (1) and the fact that $\operatorname{supp}a\cap\operatorname{supp}(1-\chi_{x})=\operatorname{supp}b\cap\operatorname{supp}(1-\chi_{\xi})=\emptyset$, we know that
@@ 2.4.2. Higher-dimensional fractal uncertainty principle
We begin with some definitions. Let $B_{r}(x)$ denote the ball centered at $x$ of radius $r$.
@@
$\mathcal{F}_{h}f(\xi)=(2\pi h)^{-\frac{n}{2}}\int_{\mathbb{R}^{n}}e^{-\frac{i}{h}\left\langle x,\xi\right\rangle}f(x)dx.$
@@
(2.20) $S(\delta)\coloneqq S+B_{\delta}(0).$
@@
(2.21) $\sup|\partial^{\alpha}\chi_{\pm}|\leq C_{\alpha}h^{-|\alpha|},$
where $C_{\alpha}$ depends on $\alpha$, $\nu$, and $n$. Therefore,
@@
Note that $\chi_{j}^{\pm}$ still satisfy the derivative bounds (2.21). With respect to the strong operator topology,
@@ 2. We will show
where $C,\beta>0$ depend only $\nu$ and $n$ and $C_{N}>0$ depends only on $N$ and $n$. Indeed, we obtain (2.24) using (2.26) for $|j-j^{\prime}|+|k-k^{\prime}|\leq h^{-\beta/2n}$ and using (2.27) with $N=\lceil 4+4n+2n/\beta\rceil$ for $|j-j^{\prime}|+|k-k^{\prime}|>h^{-\beta/2n}$. Similarly, (2.25) follows from (2.26) and (2.28).
@@
$\mathcal{K}(x,y)=(2\pi h)^{-n}\chi_{j}^{-}(x)\chi_{j^{\prime}}^{-}(y)\int_{\mathbb{R}^{n}}e^{i\left\langle y-x,\xi\right\rangle/h}\chi_{k}^{+}(\xi)\chi_{k^{\prime}}^{+}(\xi)d\xi.$
@@ 3.1. Definitions
with corresponding principal symbol
@@
The earlier works [Sch] and [dyatlov2021semiclassical], used $\Lambda=\Lambda_{+}$. However, to prove Theorem 1.3, we require more control over the precise stable/unstable behavior of $A$ on a particular subspace determined by $V$.
@@
We call elements of $\mathcal{X}$ uncontrolled long logarithmic words and elements of $\mathcal{Y}$ controlled long logarithmic words. To simplify notation, we set
@@ 3.2. Decay of $B_{\mathcal{Y}}$
$\left\langle\operatorname{Op}_{N_{j},\theta_{j}}\left(|b_{1}|^{2}\right)u_{j},u_{j}\right\rangle_{\mathcal{H}_{j}}\rightarrow\int_{\mathbb{T}^{2n}}|b_{1}|^{2}d\mu=0.$
@@
$\leq N^{-\frac{2\rho\log 2}{J\log\Lambda}}\left(\left\langle\operatorname{Op}_{N,\theta}\left(b_{d}\right)^{*}\operatorname{Op}_{N,\theta}\left(b_{d}\right)u,u\right\rangle_{\mathcal{H}}-\left\langle\operatorname{Op}_{N,\theta}\left(b_{c}\right)^{*}\operatorname{Op}_{N,\theta}\left(b_{c}\right)u,u\right\rangle_{\mathcal{H}}\right)+CN^{-\frac{4}{5}}\|u\|^{2}_{\mathcal{H}}$
@@
$\leq N^{-\frac{2\rho\log 2}{J\log\Lambda}}\left(\left\|B_{d}u\right\|^{2}_{\mathcal{H}}-\left\|B_{c}u\right\|^{2}_{\mathcal{H}}\right)+CN^{-\frac{4}{5}}\|u\|^{2}_{\mathcal{H}}.$
@@ 3.3. Decay of $B_{\mathcal{X}}$
$\mathcal{X}\leq C(\log N)^{2J}N^{\frac{2c\rho\sqrt{\alpha}}{\log\Lambda}}.$
@@
As $\mathcal{X}=\#(\mathcal{W}(T_{0})\setminus\mathcal{Z})^{2J}$, it suffices to bound $(\mathcal{W}(T_{0})\setminus\mathcal{Z})$. Recall that $\mathrm{w}=w_{0}\cdots w_{T_{0}-1}$ is in $\mathcal{W}(T_{0})\setminus\mathcal{Z}$ if and only if $\{j\in\{0,\ldots,T_{0}-1\}:w_{j}=1\}<\alpha T_{0}$. Thus, using the fact that $\alpha<1/2$ and $T_{0}=\left\lfloor\frac{\rho\log N}{J\log\Lambda}\right\rfloor$,
$\#(\mathcal{W}(T_{0})\setminus\mathcal{Z})$ $\leq\sum_{0\leq l\leq\alpha T_{0}}\binom{T_{0}}{l}\leq(\alpha T_{0}+1)\binom{T_{0}}{\left\lceil\alpha T_{0}\right\rceil}$
@@
$(\alpha\log\alpha+(1-\alpha)\log(1-\alpha))\leq c\sqrt{\alpha},$
@@ 4.1. Proof groundwork
We fix a hyperbolic $A\in\operatorname{Sp}(2n,\mathbb{Z})$ that is diagonalizable in $\mathbb{C}$. We use this diagonalizability condition in Lemma 4.1, (4.3), and (4.4). Let $M\in\mathcal{M}_{A}$ be a metaplectic transformation quantizing $A$. As $A$ is symplectic, its transpose is conjugate to its inverse. Thus, if $\lambda$ is an eigenvalue of $A$, $\lambda^{-1}$ is also an eigenvalue.
@@ 4.1.2. Construction of symbols
$\text{supp }b_{1}$$\text{supp }b_{2}$$r\sqrt{n}$$\text{supp }\tilde{b}_{1}$$r\sqrt{n}$$\text{supp }\tilde{b}_{R+1}$
Figure 1. The colored regions of (A) depict of the support of $b_{i}$ and $\tilde{b}_{i}$. Here $\Sigma^{\prime}$ is given by the black lines with $S=2$. The width of $\operatorname{supp}\tilde{b}_{i}$ is $r\sqrt{n}$. The illustration (B) demonstrates the action of $A$ on $\operatorname{supp}\tilde{b}_{R+1}$. Specifically, $A(\operatorname{supp}\tilde{b}_{R+1})$ must look like one of the two images.
@@ 4.1.3. Preliminary bound on $B_{\mathrm{w}}$
We use the notation $\mathrm{z}_{(2i-1)}\coloneqq\mathfrak{z}_{0}^{(2i-1)}\cdots\mathfrak{z}_{s_{i}-1}^{(2i-1)}$ and $\mathrm{z}_{(2i)}=\mathfrak{z}_{0}^{(2i)}\cdots\mathfrak{z}_{r_{i}-1}^{(2i)}$ and set
$\tilde{b}_{\mathrm{z}_{(2i-1)}}=\prod_{k=0}^{s_{i}-1}\tilde{b}_{\mathfrak{z}_{k}^{(2i-1)}}\circ A^{k}.$
@@
$\tilde{b}_{\mathrm{z}_{m}}=\prod_{k=0}^{2T_{1}-1}\tilde{b}_{\mathfrak{z}_{k}}\circ A^{k},$
@@
$\tilde{B}_{\mathrm{z}_{m}}\coloneqq\tilde{B}_{\mathfrak{z}_{2T_{1}-1}}(2T_{1}-1)\cdots\tilde{B}_{\mathfrak{z}_{1}}(1)\tilde{B}_{\mathfrak{z}_{0}}(0),$
@@
$\left\|\sum_{\mathrm{z}_{m}\in\mathbb{M}_{2}}\tilde{B}_{\mathrm{z}_{m}}\right\|\leq SN^{\frac{2\alpha\rho\log SR}{\log\Lambda}}\max_{\mathrm{z}_{m}\in\mathbb{M}_{2}}\|B_{\mathrm{z}_{m}}\|.$
@@
It suffices to show $\mathbb{M}_{2}\leq SN^{\frac{2\alpha\rho\log SR}{\log\Lambda}}$. We fix $1\leq i\leq K+1$ and examine $\tilde{b}_{\mathrm{z}_{(2i-1)}}$. Note that for each $1\leq k\leq S$, there exists a unique $1\leq j_{k}\leq S$ such that $A(\pi_{W^{\perp\sigma}}(x_{k}))=\pi_{W^{\perp\sigma}}(x_{j_{k}})$. In other words, $A(\operatorname{supp}\tilde{\varphi}_{R+k})\cap\operatorname{supp}\tilde{\varphi}_{R+j_{k}}\neq\emptyset$. See Figure 1. We know
@@
$\mathfrak{z}_{1}^{(2i-1)}=j_{\mathfrak{z}_{0}^{(2i-1)}},\quad\mathfrak{z}_{2}^{(2i-1)}=j_{\mathfrak{z}_{1}^{(2i-1)}},\quad\ldots,\quad\mathfrak{z}_{s_{i}-1}^{(2i-1)}=j_{\mathfrak{z}_{s_{i}-2}^{(2i-1)}}.$
Consequently, for each choice of $\mathfrak{z}_{0}^{(2i-1)}\in\{R+1,\ldots,R+S\}$, there is only one way to pick $\mathfrak{z}_{1}^{(2i-1)},\ldots,\mathfrak{z}_{s_{i}-1}^{(2i-1)}$ such that $\tilde{b}_{\mathrm{z}_{(2i-1)}}$ is nonzero. Then, there are $S$ nonzero choices of each $\tilde{b}_{\mathrm{z}_{(2i-1)}}$. Therefore, $\mathbb{M}_{2}\leq S^{K+1}R^{2\alpha T_{1}}$. By (3.3) and (4.9),
@@
We now turn our attention to $\sum_{\mathrm{z}_{m}\in\mathbb{M}_{1}}\tilde{B}_{\mathrm{z}_{m}}$.
@@
Choose $\mathrm{z}_{m}\in\mathbb{M}_{1}$. From the proof of Lemma 4.2, we see that as $\mathrm{z}_{m}\in\mathbb{M}_{1}$, there exists some $0\leq k<2T_{1}-1$ such that
$\operatorname{supp}(\tilde{b}_{\mathfrak{z}_{k}}\circ A^{k})\cap\operatorname{supp}(\tilde{b}_{\mathfrak{z}_{k+1}}\circ A^{k+1})=\emptyset.$
@@
$\|\tilde{B}_{\mathfrak{z}_{k+1}}(k+1)\tilde{B}_{\mathfrak{z}_{k}}(k)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}=\|\tilde{B}_{\mathfrak{z}_{k+1}}(1)\tilde{B}_{\mathfrak{z}_{k}}(0)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}=\mathcal{O}(N_{j}^{-\infty}).$
@@
where $h=(2\pi N)^{-1}$ and $\operatorname{Op}_{h}$ is the semiclassical Weyl quantization, defined in (2.1). Thus, using the fact that $M$ is unitary,
$\|\tilde{B}_{\mathfrak{z}_{i}}(i)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}=\|\tilde{B}_{\mathfrak{z}_{i}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\leq\|\operatorname{Op}_{h_{j}}(\tilde{b}_{\mathfrak{z}_{i}})\|_{L^{2}\rightarrow L^{2}}.$
@@
$\|\tilde{B}_{\mathrm{z}_{m}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}$ $\leq\|\tilde{B}_{\mathfrak{z}_{2T_{1}-1}}(2T_{1}-1)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\cdots\|\tilde{B}_{\mathfrak{z}_{k+1}}(k+1)\tilde{B}_{\mathfrak{z}_{k}}(k)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\cdots\|\tilde{B}_{\mathfrak{z}_{0}}(0)\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}$
@@
(4.11) $\|B_{\mathrm{w}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}\leq SN^{\frac{2\alpha\rho\log SR}{\log\Lambda}}\max_{\mathrm{z}_{m}\in\mathbb{M}_{2}}\|\tilde{B}_{\mathrm{z}_{m}}\|_{\mathcal{H}_{j}\rightarrow\mathcal{H}_{j}}+\mathcal{O}(N_{j}^{-\infty}).$
@@
Write $\mathrm{z}_{m}=\mathrm{z}_{+}\mathrm{z}_{-}$, where $\mathrm{z}_{\pm}$ are words of length $T_{1}$. Relabel $\mathrm{z}_{+}=\mathfrak{z}_{T_{1}}^{+}\cdots\mathfrak{z}_{1}^{+}$, $\mathrm{z}_{-}=\mathfrak{z}_{0}^{-}\cdots\mathfrak{z}_{T_{1}-1}^{-}$. We see that
(4.14) $\tilde{B}_{\mathrm{z}_{m}}$ $=\tilde{B}_{\mathfrak{z}^{-}_{T_{1}-1}}(2T_{1}-1)\cdots\tilde{B}_{\mathfrak{z}^{-}_{0}}(T_{1})\tilde{B}_{\mathfrak{z}_{1}^{+}}(T_{1}-1)\cdots\tilde{B}_{\mathfrak{z}_{T_{1}-1}^{+}}(0)$
@@
$\tilde{b}_{+}=\prod_{k=1}^{T_{1}}\tilde{b}_{\mathfrak{z}_{k}^{+}}\circ A^{-k}\quad\text{and}\quad\tilde{b}_{-}=\prod_{k=0}^{T_{1}-1}\tilde{b}_{\mathfrak{z}_{k}^{-}}\circ A^{k}.$
@@
We give the proof for $\tilde{b}_{-}$; the proof for $\tilde{b}_{+}$ follows similarly. By Lemma 2.8, it suffices to show for $i=1,\ldots,R+S$ and $0\leq k\leq T_{1}-1$, each $S_{L_{-},\rho,0}$-seminorm of $\tilde{b}_{i}\circ A^{k}$ is bounded uniformly in $k$ and $N$.